# Hyper-Connections

### v1

@@ 1 Introduction
[…]
@@
The two main variants of residual connections, Pre-Norm and Post-Norm, each make distinct trade-offs between gradient vanishing and representation collapse. Pre-Norm applies normalization operations to the input before each residual block, effectively addressing the problem of gradient vanishing (Bengio et al., 1994; Glorot & Bengio, 2010). However, it can also lead to the issue of collapse in deep representations (Liu et al., 2020), where hidden features in deeper layers become highly similar, diminishing the contribution of additional layers as their number increases. In contrast, Post-Norm applies normalization operations after the output of each residual block, weakening the "strength" of residuals. This approach can alleviate the issue of representation collapse but also reintroduces the problem of vanishing gradients. The vanishing gradient and the representation collapse are like two ends of a seesaw, with these two variants making respective trade-offs between these issues. The key issue is that residual connections, including both Pre-Norm and Post-Norm variants, predefine the strength of connections between the output and input within a layer.
Figure 2: Hyper-connections with expansion rate $n=2$. (a) The highlighted green and blue connections compose depth-connections between the output of the layer and the blue hidden vector. (b) The highlighted blue and yellow connections represent a part of width-connections between the hidden vectors $h_{1}$ (blue) and $h_{2}$ (yellow). (c) The matrix representation of hyper-connections.
@@
The core idea of hyper-connections is to propose learnable depth-connections and width-connections, as depicted in Fig. 2. Depth-connections can be considered as a generalized form of residual connections, assigning weights to the connections between the inputs and outputs of each layer. To enable the network to model different depth-connections simultaneously, we expand the network’s input into $n$ copies, each having its own depth connection, as shown in Fig. 2 (a). Moreover, we establish width connections between the $n$ hidden vectors, allowing information exchange between hidden vectors within the same layer, as shown in Fig. 2 (b). Finally, we can formalize the hyper-connections into a matrix, as shown in Fig. 2 (c). We find that hyper-connections can not only learn to adjust the strength of residuals but also learn to rearrange the layers (sequential or parallel), as discussed in § 3.2. Furthermore, we introduce dynamic hyper-connections (DHC), enabling the network to adjust connection weights according to the input. Notably, although hyper-connections seem to increase the network’s width by $n$ times, the additional parameters and computational cost are almost negligible.
Our research, primarily centered on large language models (LLMs) pre-training, also extends to visual generation and classification tasks. Using Pre-Norm as a baseline, we demonstrate the significant benefits of hyper-connections, including their effectiveness in 1B and 7B dense models as well as MoE models with 7B parameters, as detailed in § 5. The benefits are particularly prominent for OLMoE (Muennighoff et al., 2024) as presented in Fig.1. The model utilizing DHC converges 1.8 times faster and shows an improvement of approximately 6 points on ARC-Challenge compared to the baseline trained with 500 B tokens. According to our visual analysis, as shown in Fig.3, the baseline model tends toward representation collapse, characterized by high similarity between features of adjacent layers. In contrast, models with hyper-connections exhibit significantly lower similarity between features across adjacent layers and a wider range of similarities. This suggests that hyper-connections enhance the impact of each layer. Further discussion is provided in §4.5 and in Appendix D. These compelling pieces of evidence demonstrate the generality of the hyper-connections principle, and we anticipate their applicability in numerous other AI challenges.
@@ 2.1 Static Hyper-Connections
Consider the hidden vector $\mathbf{h}^{k-1}\in\mathbb{R}^{d}$ (or $\mathbf{h}^{k-1}\in\mathbb{R}^{d\times 1}$) as the input to the $k$-th layer, with the initial input $\mathbf{h}^{0}$ to the network. Initially, $\mathbf{h}^{0}\in\mathbb{R}^{d}$ is replicated $n$ times to form the initial hyper hidden matrix $\mathbf{H}^{0}=\begin{pmatrix}\mathbf{h}^{0}&\mathbf{h}^{0}&\dots&\mathbf{h}^{0}\end{pmatrix}^{\intercal}\in\mathbb{R}^{n\times d}$. Here, $n$ is called the expansion rate. For the $k$-th layer, the input consists of the hyper hidden matrix from the previous layer $\mathbf{H}^{k-1}=\begin{pmatrix}\mathbf{h}^{k-1}_{1}&\mathbf{h}^{k-1}_{2}&\dots&\mathbf{h}^{k-1}_{n}\end{pmatrix}^{\intercal}\in\mathbb{R}^{n\times d}$. Finally, we sum the last hyper hidden matrix row-wise to obtain the required hidden vector, which is then passed through a final projector to produce the final output of the network (i.e., a normalization layer and an unembedding layer in transformers). To simplify the notation in subsequent analysis, we omit the layer index and simply denote the hyper-hidden matrix as $\mathbf{H}=\begin{pmatrix}\mathbf{h}_{1}&\mathbf{h}_{2}&\dots&\mathbf{h}_{n}\end{pmatrix}^{\intercal}$.
The hyper-connections can be represented by a matrix $\mathcal{HC}$, where each element defines the connection weight. The matrix is structured as follows:
@@
Consider a network layer, denoted by $\mathcal{T}$, it integrates self-attention layers and feed-forward networks within transformers. The output of the hyper-connections, denoted by $\mathbf{\hat{H}}$, can be simply formulated as follows:
@@
While $\mathbf{A_{r}}$ is used to connect $\mathbf{H}$ and map it into a residual hyper hidden matrix $\mathbf{H^{\prime}}$. It is represented as follows:
@@ 2.2 Dynamic Hyper-Connections
Our experiments in § 5 demonstrate that dynamic hyper-connections outperform static hyper-connections in language modeling tasks. The PyTorch implementations for both the static and dynamic variants of hyper-connections are detailed in Algorithm 2.
@@ 3.1 Residual Connections as Non-trainable Hyper-Connections
For Pre-Norm, its hyper-connection matrix is a $2\times 2$ matrix where the bottom right triangular part is filled with 11 and the rest is a placeholder 0. For Post-Norm, the weights depend on the variances and covariance of the input and output, forming a $2\times 2$ matrix. Therefore, their hyper-connection matrices are non-trainable. In this work, we propose hyper-connections that can be $(n+1)\times(n+1)$ matrices, with weights that are trainable or even predicted based on the input. The complete derivation is provided in Appendix E.
@@ 3.2 Sequential-Parallel Duality
When the hyper-connections for odd and even layers (with layer numbering starting from 1) are defined by the following matrices, the neural network will be arranged in parallel every two consecutive layers, similar to the arrangement of parallel transformer blocks in transformers (Wang, 2021), as shown in Fig. 4 (b). The general and complete derivation is provided in Appendix F.
@@ 4 Results
Figure 5: Comparison of training loss curves for different expansion rate. The left subfigure includes models with dynamic hyper-connections (DHC) at various expansion rates, while the right subfigure shows the effect of omitting the tanh function. Both subfigures illustrate how increasing the expansion rate leads to improved training loss performance over 500500B tokens. Results are smoothed using an exponential moving average with a coefficient of 0.99.
@@ Methods
[…]
Table 1: Performance comparison of different expansion rates on 500 B tokens.
We primarily conduct experiments on pre-training of large language model, including dense and Mixture-of-Experts (MoE) (Shazeer et al., 2017) models, and extend to visual generation and classification tasks. Due to space constraints, we include the vision experiments in the Appendix C.
Experiment Settings. We employ the experimental setup outlined by OLMo (Groeneveld et al., 2024) for dense models and by OLMoE (Muennighoff et al., 2024) for MoE models. For dense models, we use dolmap-v1.5-sample (Soldaini et al., 2024) as our training dataset. We conduct ablation studies on 11B models and assess the effectiveness of our method at the 77B model scale. For MoE models, we train the OLMoE-1B-7B model, both with and without hyper-connections, on the OLMOE-MIX dataset. These models activate 1.31.3B out of a total of 77B parameters. All experiments are trained on 500500B tokens.
Implementation. We maintain the training configuration of the baseline model, replacing the residual connections with hyper-connections. The static component in Eqs. 1, 11, 12, 13 does not utilize weight decay, whereas the dynamic component does. Since the hyper hidden vectors of the final transformer block are ultimately summed, we ensure that the standard deviation (std) of the output (before the final layernorm and unembedding layers) remains consistent with the original. At initialization, we scale the std of the weights of the output module at all layers, including those of the second linear layer of the feedforward network and the output projector of the attention module, by a factor of $\sqrt{n}$, where $n$ represents the expansion rate.
@@ 4.1 Ablation Study
[…]
Staitc and dynamic hyper-connections. Table 2 presents an ablation study comparing SHC and DHC. All hyper-connection (HC) variants significantly outperform the baseline. At an expansion rate of 2, the improvements of DHC and SHC are similar. However, at an expansion rate of 4, DHC performs notably better than SHC.
@@ Methods
[…]
Table 2: Ablation Study on Static and Dynamic Hyper-Connections with Training on 500 B Tokens
@@
[…]
Table 4: Performance of Related Methods on OLMo-1B.
We implemented the Altup (Baykal et al., 2024) and ResiDual (Xie et al., 2023) methods in OLMo. Altup is motivated to widen the hidden dimension while maintaining low computation cost by passing only a part of hidden state to transformer blocks. By contrast, ResiDual is proposed to combine both Pre- and Post-Norm in a two-stream style. Both methods expand the hidden size by $n$ times with negligible computational overhead, with ResiDual expanding it exactly $2$ times. For a fair comparison, we set $n=2$ in our experiments. Unfortunately, these methods show gains in the early stages of training but are gradually surpassed by the baseline, as shown in Table 4.
@@
[…]
@@
Down Stream
Avg, Acc. $\uparrow$
OLMo-7B 2.581 14.316 2.322 11.324 70.1
[…]
Table 5: Evaluation Metrics for 7B Models
[…]
@@ 4.5 Visualization Analysis
where $c_{kj}^{(0)}$ describes how much layer-$j$ ($\mathcal{T}^{j}$) contributes to layer-$k$’s input $\mathbf{h}_{0}^{k}$. Then, $\mathbf{C}^{(0)}$ denotes a dense connection weight matrix. In particular, let layer-0 be the word embedding and $\mathcal{T}^{0}$ be an identity mapping, layer-$L+1$ be the hidden state before the unembedding layer, which is a summation over the last hidden vectors, i.e., $\mathbf{h}_{0}^{L+1}=\sum_{j}\mathbf{h}_{j}^{L}$.
[…]

### v2

@@ 1 Introduction
Figure 1: The performance of the baseline model OLMoE-1B-7B and the model with hyper-connections, OLMoE-1B-7B-DHC×$\times$4. (1) and (2) show the training loss (0.99 EMA smoothed) and the C4-en validation loss, respectively. Our method converges 1.8 times faster compared to the baseline and maintains a significant advantage at the 500B tokens. (3) and (4) show the accuracy curves on HellaSwag and ARC-Challenge, demonstrating the superior performance of the OLMoE-1B-7B-DHC×$\times$4 model.
@@
The two main variants of residual connections, Pre-Norm and Post-Norm, each make distinct trade-offs between gradient vanishing and representation collapse. Pre-Norm applies normalization operations to the input before each residual block, effectively addressing the problem of gradient vanishing (Bengio et al., 1994; Glorot & Bengio, 2010). However, it can also lead to the issue of collapse in deep representations (Liu et al., 2020), where hidden features in deeper layers become highly similar, diminishing the contribution of additional layers as their number increases. In contrast, Post-Norm applies normalization after the output of each residual block, reducing the influence of a hidden state on subsequent layers. This approach can alleviate the issue of representation collapse but also reintroduces the problem of vanishing gradients. The vanishing gradient and the representation collapse are like two ends of a seesaw, with these two variants making respective trade-offs between these issues. The key issue is that residual connections, including both Pre-Norm and Post-Norm variants, predefine the strength of connections between the output and input within a layer.
Figure 2: Hyper-connections (HC) with an expansion rate of $n=2$. (a) Residual connections. (b) Hyper-connections: $\beta_{1}$, $\beta_{2}$, $\alpha_{0,0}$, $\alpha_{0,1}$, $\alpha_{1,0}$, $\alpha_{1,1}$, $\alpha_{2,1}$, and $\alpha_{2,2}$ are learnable scalars or scalars predicted by the network , depending on the specific HC version. These connections enable lateral information exchange and vertical integration of features across depths. The Transformer with HC is shown in Fig. 17. They can be decoupled into depth-connections and width-connections. (c) Depth-connections perform a weighted sum between the layer output and the hidden vector $h_{1}$. (d) Width-connections allow information exchange between the hidden vectors $h_{1}$ and $h_{2}$.
@@
The core idea of hyper-connections (HC) is to propose learnable depth-connections and width-connections, as depicted in Fig.2 (b). These connections flexibly integrate features vertically across depths, compared to the residual connections shown in Fig.2 (a). Depth-connections can be considered as a generalized residual connections, assigning weights to the connections between the inputs and outputs of each layer. To enable the network to model different depth-connections simultaneously, we expand the network’s input into $n$ copies, each having its own depth connection, as shown in Fig. 2 (b). This design allows multiple hidden vectors to reserve multiple patterns connecting preceding layers, as shown in § 4.5. Moreover, we establish width connections between the $n$ hidden vectors, allowing information exchange between hidden vectors within the same layer, as shown in Fig. 2 (b). We argue that $n$ ($>1$) hidden states are necessary. As analyzed in Appendix F, the seesaw effect persists when $n=1$, and experiments show that it does not improve performance, as shown in Fig. 5. In contrast, when $n>1$, hyper-connections can not only learn to adjust the strength of residuals but also rearrange layers, either sequentially or in parallel, as discussed in § 3.2. To further enhance flexibility, we introduce dynamic hyper-connections (DHC), enabling the network to adjust connection weights according to the input. Notably, although HC seem to increase the network’s width by $n$ times, the additional parameters and computational cost are almost negligible, as analyzed in Appendix B. The Transformer with HC is shown in Fig. 17.
Our research, primarily centered on large language models (LLMs) pre-training, also extends to visual generation and classification tasks. Using Pre-Norm as a baseline, we demonstrate the significant benefits of hyper-connections, including 1B and 7B dense models as well as 7B MoE models, as detailed in § 5. The benefits are particularly prominent for OLMoE (Muennighoff et al., 2024) as presented in Fig.1. The model utilizing DHC converges 1.8 times faster and shows an improvement of 6 points on ARC-Challenge compared to the baseline trained with 500 B tokens. According to our visualization analysis, as shown in Fig.3, the baseline model tends toward representation collapse, characterized by high similarity between features of adjacent layers. In contrast, models with HC exhibit significantly lower similarity between features across adjacent layers and a wider range of similarities. This suggests that HC enhance the impact of each layer. Further discussion is provided in §4.5 and in Appendix F. These compelling pieces of evidence demonstrate the generality of the hyper-connections principle, and we anticipate their applicability in numerous other AI challenges.
@@ 2.1 Static Hyper-Connections
Consider the hidden vector $\mathbf{h}^{k-1}\in\mathbb{R}^{d}$ (or $\mathbf{h}^{k-1}\in\mathbb{R}^{d\times 1}$) as the input to the $k$-th layer, with the initial input $\mathbf{h}^{0}$ to the network. Initially, $\mathbf{h}^{0}\in\mathbb{R}^{d}$ is replicated $n$ times to form the initial hyper hidden matrix $\mathbf{H}^{0}=\begin{pmatrix}\mathbf{h}^{0}&\mathbf{h}^{0}&\dots&\mathbf{h}^{0}\end{pmatrix}^{\intercal}\in\mathbb{R}^{n\times d}$. Here, $n$ is the expansion rate. For the $k$-th layer, the input consists of the hyper hidden matrix from the previous layer $\mathbf{H}^{k-1}=\begin{pmatrix}\mathbf{h}^{k-1}_{1}&\mathbf{h}^{k-1}_{2}&\dots&\mathbf{h}^{k-1}_{n}\end{pmatrix}^{\intercal}\in\mathbb{R}^{n\times d}$. Finally, we sum the last hyper hidden matrix row-wise to obtain the required hidden vector, which is then passed through a final projector to produce the final output of the network (i.e., a normalization layer and an unembedding layer in transformers). To simplify the notation in subsequent analysis, we omit the layer index and simply denote the hyper-hidden matrix as $\mathbf{H}=\begin{pmatrix}\mathbf{h}_{1}&\mathbf{h}_{2}&\dots&\mathbf{h}_{n}\end{pmatrix}^{\intercal}$.
The hyper-connections (HC) can be represented by a matrix $\mathcal{HC}$, where each element defines the connection weight. The matrix is structured as follows:
@@
Consider a network layer $\mathcal{T}$, it integrates self-attention layers and feed-forward networks within transformers. The output of the HC, denoted by $\mathbf{\hat{H}}$, can be simply formulated as follows:
@@
While $\mathbf{A_{r}}$ is used to connect $\mathbf{H}$ and map it to a hyper hidden matrix $\mathbf{H^{\prime}}$, as shown below:
@@ 2.2 Dynamic Hyper-Connections
Our experiments in § 5 demonstrate that dynamic hyper-connections outperform static hyper-connections in language modeling tasks. The PyTorch implementations for both the static and dynamic variants of hyper-connections are detailed in Algorithm 2 and 3.
@@ 3.1 Residual Connections as Non-trainable Hyper-Connections
For Pre-Norm, its hyper-connection matrix is a $2\times 2$ matrix where the bottom right triangular part is filled with $1$ and the rest is a placeholder 00. For Post-Norm, the weights depend on the variances and covariance of the input and output, forming a $2\times 2$ matrix. Therefore, their hyper-connection matrices are non-trainable. In this work, we propose hyper-connections that can be $(n+1)\times(n+1)$ matrices, with weights that are trainable or even predicted based on the input. The complete derivation is provided in Appendix G.
@@ 3.2 Sequential-Parallel Duality
When the hyper-connections for odd and even layers (with layer numbering starting from 1) are defined by the following matrices, the neural network will be arranged in parallel every two consecutive layers, similar to the arrangement of parallel transformer blocks in transformers (Wang, 2021), as shown in Fig. 4 (b). The general and complete derivation is provided in Appendix H.
@@ 4 Results
Figure 5: Comparison of training loss curves for different expansion rate. The left subfigure includes models with dynamic hyper-connections (DHC) at various expansion rates, while the right subfigure shows the effect of omitting the tanh function. Both subfigures illustrate how increasing the expansion rate leads to improved training loss performance over $500$B tokens. Results are smoothed using an exponential moving average with a coefficient of 0.99.
Table 1: Ablation study on expansion rates $n$ with training on 500 B tokens.
@@ Methods
OLMo-1B-DHC×$\times$1 W/O tanh 2.822 18.270 2.556 14.428 62.3
OLMo-1B-DHC×$\times$2 W/O tanh 2.792 17.663 2.537 14.033 63.8
OLMo-1B-DHC×$\times$4 W/O tanh 2.779 17.451 2.516 13.844 64.4
OLMo-1B-DHC×$\times$8 W/O tanh 2.777 17.425 2.514 13.819 63.8
OLMo-1B-DHC×$\times$1 2.819 18.125 2.556 14.418 62.3
OLMo-1B-DHC×$\times$2 2.802 17.950 2.534 14.114 63.0
OLMo-1B-DHC×$\times$4 2.781 17.509 2.514 13.826 63.8
OLMo-1B-DHC×$\times$8 2.778 17.445 2.516 13.843 62.8
We primarily conduct experiments on pre-training of large language model, including dense and Mixture-of-Experts (MoE) (Shazeer et al., 2017) models, and extend to visual generation and classification tasks. Due to space constraints, we include the vision experiments in the Appendix E.
Experiment Settings. We employ the experimental setup outlined by OLMo (Groeneveld et al., 2024) for dense models and by OLMoE (Muennighoff et al., 2024) for MoE models. For dense models, we use dolmap-v1.5-sample (Soldaini et al., 2024) as our training dataset. We conduct ablation studies on $1$B models and assess the effectiveness of our method at the $7$B model scale. For MoE models, we train the OLMoE-1B-7B model, both with and without hyper-connections, on the OLMOE-MIX dataset. These models activate $1.3$B out of a total of $7$B parameters. All experiments are trained on $500$B tokens.
Implementation. We maintain the training configuration of the baseline model, replacing the residual connections with hyper-connections. The static component in Eqs. 1, 11, 12, 13 does not utilize weight decay, whereas the dynamic component does. Since the hyper hidden vectors of the final transformer block are ultimately summed, we ensure that the standard deviation (std) of the output (before the final layernorm and unembedding layers) remains consistent with the original. At initialization, we scale the std of the weights of the output module at all layers, including those of the second linear layer of the feedforward network and the output projector of the attention module, by a factor of $\sqrt{n}$, where $n$ represents the expansion rate. The parameters and computational overhead introduced by hyper-connections is negligible, see Table. 7 and  8.
@@ 4.1 Ablation Study
The evaluation results are presented in Table 1, and the training loss curves are depicted in Fig. 5. We observe that with an expansion rate of $n=1$, the performance of DHC is inferior to the baseline. However, for $n>1$, DHC significantly outperforms the baseline, achieving superior results at $n=4$, with the increase to $n=8$ providing minimal additional benefits. Notably, OLMo-1B-DHC×$\times$8 W/O tanh excels on both V2 and V3 validation sets, with a reduction in V2 Eval Loss by 0.034 and V3 Eval Loss by 0.029 compared to the baseline. Furthermore, the decline rate of training losses for DHC ($n\geq 2$) is steeper than that of the baseline, and DHC demonstrates greater stability, with no spikes observed in any DHC experiments.
Static and dynamic hyper-connections. Table 2 presents an ablation study comparing SHC and DHC. All hyper-connection (HC) variants significantly outperform the baseline. At an expansion rate of 2, the improvements of DHC and SHC are similar. However, at an expansion rate of 4, DHC performs notably better than SHC.
Table 2: Ablation study on static and dynamic hyper-connections with training on 500 B tokens.
@@ Methods
OLMo-1B-SHC×$\times$2 2.799 17.778 2.538 14.152 63.4
OLMo-1B-DHC×$\times$2 2.802 17.950 2.534 14.114 63.0
OLMo-1B-DHC×$\times$2 W/O tanh 2.792 17.663 2.529 14.033 63.8
OLMo-1B-SHC×$\times$4 2.791 17.671 2.528 14.025 63.6
OLMo-1B-DHC×$\times$4 2.781 17.509 2.515 13.826 63.8
OLMo-1B-DHC×$\times$4 W/O tanh 2.779 17.451 2.516 13.844 64.4
@@
Table 3: Ablation study on OLMo-1B-DHC×$\times$4. In the $\mathbf{B}$ or $\mathcal{WC}$ column, the symbol "✗" denotes parameters that are not trainable from initialization.
@@ 4.2 Comparison with related works
Table 4: Performance of related methods on OLMo-1B models.
@@ Methods
OLMo-1B-Altup×$\times$2 2.827 18.268 2.558 14.454 62.4
OLMo-1B-DHC×$\times$2 2.802 17.950 2.534 14.114 63.0
OLMo-1B-DHC×$\times$2 W/O tanh 2.792 17.663 2.529 14.033 63.8
We implemented the Altup (Baykal et al., 2024) and ResiDual (Xie et al., 2023) methods in OLMo. Altup is motivated to widen the hidden dimension while maintaining low computation cost by passing only a part of hidden state to transformer blocks. By contrast, ResiDual is proposed to combine both Pre- and Post-Norm in a two-stream style. Both methods expand the hidden size by $n$ times with negligible computational overhead, with ResiDual expanding it exactly $2$ times. For a fair comparison, we set $n=2$ in our experiments. Unfortunately, although these methods show gains in the early stages of training, they are gradually surpassed by the baseline, as demonstrated by the results in Table 4 and the training loss curves in Fig. 15.
@@
Figure 6: (1) and (2) Training loss (0.99 EMA smoothed) and C4-en validation loss for OLMo-7B and OLMo-7B-DHC×$\times$4 models. (3) and (4) Accuracy curves on hellaswag and sciq, demonstrating the superior performance of the OLMo-7B-DHC×$\times$4 model.
We evaluate the effectiveness of hyper-connections on the 7B model, training a model with DHCs with an expansion rate of 4, denoted as OLMo-7B-DHC×$\times$4. According to Table 5, OLMo-7B-DHC×$\times$4 significantly outperforms the baseline OLMo-7B model in all average metrics. In the V2 evaluation, OLMo-7B-DHC×$\times$4 shows improvements of 0.022 for loss and 0.293 for PPL. Furthermore, the average score of downstream benchmarks 0.710 surpasses the baseline 0.701, with the results of specific tasks shown in Fig. 10.
Table 5: Performance of 7B models. FLOPs refers to the computation per token in the forward pass.
@@
Tasks Avg.
Acc. $\uparrow$
OLMo-7B 6.9 13.36 2.581 14.316 2.322 11.324 70.1
OLMo-7B-DHC×$\times$4 6.9 13.38 2.559 14.023 2.304 11.120 71.0
Based on Fig 6, the OLMo-7B-DHC×$\times$4 model consistently shows better metrics compared to baseline, including training and validation loss and accuracy in downstream benchmarks. Notably, after 400 B tokens, the model maintains its improvement without the gains diminishing. This indicates that the OLMo-7B-DHC×$\times$4 model continues to provide consistent benefits in reducing loss, even at higher token counts. Furthermore, according to Fig. 6, the baseline model exhibits frequent spikes, while our model with DHCs shows no spikes throughout the training. This shows that our approach not only achieves better loss but also ensures more stable training.
@@ 4.5 Visualization Analysis
where $c_{kj}^{(0)}$ describes how much layer-$j$ ($\mathcal{T}^{j}$) contributes to layer-$k$’s input $\mathbf{h}_{0}^{k}$. Then, $\mathbf{C}^{(0)}$ denotes a dense connection weight matrix. In particular, let layer-00 be the word embedding and $\mathcal{T}^{0}$ be an identity mapping, layer-$L+1$ be the hidden state before the unembedding layer, which is a summation over the last hidden vectors, i.e., $\mathbf{h}_{0}^{L+1}=\sum_{j}\mathbf{h}_{j}^{L}$.
OLMo-1B-DHC×$\times$4 model is adopted for visualization. We take the checkpoint at 500B tokens and forward random validation text to obtain dynamic hyper-connection weights. In addition, we show connection patterns for some related baseline methods. Finally, the visualization is illustrated in Fig. 13. We present the following findings, with more detailed discussions provided in Appendix F.